VEGAS Team Meeting

March 25, 2013


CALTECH, UC-B, & Columbia

Dan, Andrew, Jayanth, Glenn, Mark


Jason, Ray, Marty, John, Anish, Paul, Mark, Joe Brandt


  • GC Observation Preparedness
  • Ongoing system issues
  • Mode 13 packet loss

Meeting Notes:

  1. are there more tests or data analysis we should do before enabling the galactic center observations?
  • ready to go - if team (Anish, Ray, & Glenn) are available to support the obs.
  • test a bright pulsar as the first measurement (~2 hrs) with the Ku Wideband receiver Not critical LST
  • A long GC observing slot would be useful so no interruption of the track
  1. what wierd stuff has been happening in our tests, and why, and do we need to do anything about it?
  • crashing halfway through the run?
  • is the frequency flipping OK?
  1. jayanth and john's testing of mode13 packet loss. paul suggested we turn off GUPPI. what else?
  • GUPPI is off now and will test today.

Next Teleconference: April 01, 2013 .

Latest test results

Long-term action Items
  1. Ray to make VEGAS simulator.
  2. A BOF for mode 13, or some other mode that uses GPUs.
    • (04-Feb) Not available yet
    • (11-Feb) 16 input version compiling now for subsequent test
    • (04-Mar) Should be ready to test in GB ~7-Mar
      • Need to refine what tests are done in Berkeley v. GB
        • Noise + tone tests in Berkeley
        • Replicate in GB (one Roach)
        • Build-out to eight Roach boards
        • Test on GBT
  3. A BOF with working switching signal logic.
    • (04-Feb) designed but still troubleshooting switching table loading issues
    • (11-Feb) Added to design for Mode #1
    • (11-Feb) Add sfp shift into manager
  4. A test and decision on what computers to buy.
    • (04-Feb) Not yet determined - Jayanth should go ahead with his tests to complete selection
    • (04-Feb) Needs HPC11.
  5. All Banks not starting at the same time.
    • (04-Feb) (above)
    • DONE (04-Mar) Resolved
  6. It will be good if we get the full spectral dynamic range corresponding to the 8 bit ADC in the next version of mode 1.
    • (04-Feb) Conduct initial tests at Berkeley, to see if there are any limiting issues for dynamic range and optimize
    • (04-Feb) what is/are the requirement(s) w.r.t full scale sin wave?
    • (04-Feb) is this a new requirement for the manager (yes)? is a parameter required to select between options? (parametrized shifting)
    • (04-Feb) what are requirements for shift parameter? depends on result of studying spectra...
    • (25-Feb)
  7. KFPA observation with mode 13.
    • (04-Feb) awaits mode 13
  8. Manager/VEGAS HPC code that will let two instances of the program be run simultaneously.
    • (04-Feb) not available, but not required for GC obs.
  9. Remember to do the RFI test of the rack.
    • (04-Feb) Not yet scheduled
  10. Roach to Valon serial port connection.
    • (04-Feb) Not yet implemented; not req'd for GC obs
  11. The I2C bus software tested with the manager and Galen's modules.
    • (04-Feb) Post GC Obs
  12. Add software version information into FITS file
    • (11-Feb) After GC obs

Action Items:

Other items:

-- MartyBloss - 2013-03-25
Topic revision: r1 - 2013-03-25, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback